The sum of three numbers is 119. The second number is 4 times the third. The third number is 7 more than the first. What are the numbers?
Answer by Edwin McCravy(20056)

Let x = the first number
The third number is 7 more than the first.
So the third number = x+7
The second number is 4 times the third.
So the second number = 4(x+7)
The sum of three numbers is 119.
So x + x+7 + 4(x+7) = 118. 

Solve that for x.  It will be the first number
Add 7 to get the third number.  Then multiply
that by 4 to get the second number.

Edwin